OpenStudy (anonymous):

How many liters of chlorine gas, at standard temperature and pressure, will react with 15.5 grams of potassium metal? Show all of the work used to solve this problem. 2 K + Cl2 yields 2 KCl

OpenStudy (paxpolaris):

since, the molar mass of Potassium (K) is 39.1 g:\[15.5\ \cancel{g\ K} \times {1\ mol\ K \over 39.1\ \cancel{g\ K}}= \color{green}{0.396\ mol\ K}\]

OpenStudy (paxpolaris):

and as we need 2 moles of K, for every 1 mole of Cl2 gas ... (from the equation) , we will only need half as many moles of Cl2\[0.396 \times \frac12=\color{green}{0.198\ mol\ Cl_2}\]

OpenStudy (paxpolaris):

Finally, as gases have cool property that at Standard Temperature and Pressure (STP): \(1\) mole of any gas will occupy exactly \(22.4\) liters of space. \[0.198\ mol\ Cl_2 \times {22. 4\ liters\ Cl_2 \over 1\ mol\ Cl_2} = \Large \color{green}{4.44\ liters\ Cl_2}\]

OpenStudy (paxpolaris):

if you want to do it all in one step using Dimensional Analysis (Unit Analysis):\[15.5\ \color{maroon}{g\ K} \times {1\ \color{navy}{mol\ K} \over 39.1\ \color{maroon}{g\ K}}\times {1\ \color{olive} {mol\ Cl_2} \over 2\ \color{navy}{mol\ K}}\times {22. 4\ \color{orange}{liters\ Cl_2} \over 1\ \color{olive} {mol\ Cl_2}}=\LARGE 4.44\ \color{orange}{liters\ Cl_2}\]

OpenStudy (anonymous):

@PaxPolaris how do I find the molar mass of Potassium?

OpenStudy (aaronq):

the molar mass of potassium (and all other elements) can be found on the periodic table of elements.

OpenStudy (anonymous):

thank you